The revenue function for a school group selling n bookmarks is given by R(n) = 2n, and the total cost function is given by C(n) = 144+0.8n. Determine the number of bookmarks that need to be sold for the school group to break even
Answer by stanbon(75887)

The revenue function for a school group selling n bookmarks is given by R(n) = 2n, and the total cost function is given by C(n) = 144+0.8n. Determine the number of bookmarks that need to be sold for the school group to break even
-----
Break even implies revenue = cost
------
2n = 144+0.8n
------
1.2n = 144
n = 144/1.2 = 1440/12 = 120
-------------
Cheers,
Stan H.
---------------